### v1

# The most symmetric smooth cubic surface over a finite field of characteristic $2$

Abstract.
In this paper we find the largest automorphism group of a smooth cubic surface over any finite field of characteristic $2$ We prove that if the order of the field is a power of $4,$ then the automorphism group of maximal order of a smooth cubic surface over this field is $\mathrm{PSU}_{4}(\mathbb{F}_{2}).$ If the order of the field of characteristic $2$ is not a power of $4,$ then we prove that the automorphism group of maximal order of a smooth cubic surface over this field is the symmetric group of degree $6$ Moreover, we prove that smooth cubic surfaces with such properties are unique up to isomorphism.
1. Introduction
The Fermat cubic surface
$x^{3}+y^{3}+z^{3}+t^{3}=0$
was interesting to many mathematicians since ancient times, at least since Porisms of Diophantus, where he mentioned that the difference of two cubes can be represented as a sum of two cubes (see [4, Chapter \@slowromancapxxi@] for more historical details). It was studied a lot in number theory and algebraic geometry. The Fermat cubic surface has the largest automorphism group among smooth cubic surfaces over an algebraically closed field of characteristic not equal to $3$ This was proved by T. Hosoh in [8, Theorem 5.3] for fields of characteristic zero and, by I. Dolgachev and A. Duncan in [6, Theorem 1.1] for fields of positive characteristics. In particular, we have the following theorem:
Theorem 1.1 ([6, Theorem 1.1]).
[…]
Note that from the proof of Theorem 1.1 (see [6, §5.1] for more details) we get the following theorem about automorphism group of the Fermat cubic surface over $\mathbb{F}_{4}.$
Theorem 1.2 ([6, §5.1]).
Let $S$ be the Fermat cubic surface over $\mathbb{F}_{4}.$ Then we have
$\mathrm{Aut}(S)\simeq\mathrm{PSU}_{4}(\mathbb{F}_{2}).$
The automorphism group of a smooth cubic surface over a finite field can be much smaller than the automorphism group of a smooth cubic surface over an algebraic closure of the field. For example, there is the following result.
Theorem 1.3 ([13, Theorem 3 and Example 9]).
Let $S$ be a smooth cubic surface over the field $\mathbb{F}_{2}$ of $2$ elements. Then $|\mathrm{Aut}(S)|\leqslant 720.$ Moreover, $|\mathrm{Aut}(S)|=720$ if and only if $\mathrm{Aut}(S)\simeq\mathfrak{S}_{6}$ and $S$ is isomorphic to the cubic surface defined by the equation
(1.1) $x^{2}t+y^{2}z+z^{2}y+t^{2}x=0.$
In this paper we find automorphism groups of smooth cubic surfaces of maximal order over all finite fields $\mathbb{F}_{2^{k}}$ of order $2^{k}$ for all natural $k.$ So we prove the following theorem.
Theorem 1.4.
Let $S$ be a smooth cubic surface over a finite field $\mathbf{F}$ of characteristic $2$
  1. (i)
[…]
    $\mathrm{Aut}(S)\simeq\mathrm{PSU}_{4}(\mathbb{F}_{2})$
    and $S$ is isomorphic to the Fermat cubic surface.
  2. (ii)
    If $\mathbf{F}=\mathbb{F}_{2^{2k+1}},$ then $|\mathrm{Aut}(S)|\leqslant 720.$ Moreover, $|\mathrm{Aut}(S)|=720$ if and only if
    $\mathrm{Aut}(S)\simeq\mathfrak{S}_{6}$
    and $S$ is isomorphic to cubic surface (1.1).
In the process of proving Theorem 1.4 we get the following result.
Theorem 1.5.
Let $S$ be a smooth cubic surface over a finite field $\mathbf{F}$ of characteristic $2$ Denote by $\Gamma$ the image of the Galois group $\mathrm{Gal}(\overline{\mathbf{F}}/\mathbf{F})$ in the Weyl group $W(\mathrm{E}_{6})\simeq\mathrm{Aut}(\mathrm{Pic}(S_{\overline{\mathbf{F}}})).$ Assume that
$\mathbb{Z}/5\mathbb{Z}\subset\mathrm{Aut}(S).$
  1. (i)
    If $\mathbf{F}=\mathbb{F}_{4^{k}}$ and $\Gamma=\{e\},$ then $S$ is isomorphic to the Fermat cubic surface.
  2. (ii)
    If $\mathbf{F}=\mathbb{F}_{2^{2k+1}}$ and $\Gamma\simeq\mathbb{Z}/2\mathbb{Z},$ then $S$ is isomorphic to cubic surface (1.1).
The following proposition shows that over fields $\mathbb{F}_{2^{2k+1}}$ the Fermat cubic surface and the cubic (1.1) are not isomorphic whereas over $\overline{\mathbb{F}}_{2}$ they are isomorphic.
Proposition 1.6 (cf. Theorem 1.4(i)).
Let $\mathcal{F}$ be the Fermat cubic surface over a field $\mathbf{F}$ of characteristic $2$ Then if $\mathbf{F}=\mathbb{F}_{2^{2k+1}}$ we get $\mathrm{Aut}(\mathcal{F})\simeq\mathbb{Z}/2\mathbb{Z}\times\mathfrak{S}_{4}.$
Corollary 1.7.
Let $\mathbf{F}$ be a field of characteristic $2$ Then
  1. (i)
    if $\mathbf{F}=\mathbb{F}_{2^{2k+1}},$ then smooth cubic surface (1.1) is not isomorphic to the Fermat cubic surface;
  2. (ii)
    if $\mathbf{F}=\mathbb{F}_{4^{k}}$ or $\mathbf{F}=\overline{\mathbf{F}},$ then smooth cubic surface (1.1) and the Fermat cubic surface are isomorphic.
Let us briefly describe the idea of the proof of Theorem 1.4. We show that we can $\mathbb{Z}/5\mathbb{Z}$-equivariantly blow down $5$ skew lines all of which are defined over the field such that we get either $\mathbb{P}^{1}\times\mathbb{P}^{1},$ or a quadric isomorphic to the Weil restriction of scalars of $\mathbb{P}^{1}.$ In the former case the rank of the Picard group of the blowdown is equal to $2,$ while in the latter case the rank of the Picard group of the blowdown is equal to $1$ So we get $5$ points in general position. We show that such set of $5$ points on a quadric is unique up to automorphisms of the quadric. So from the uniqueness of $5$ points in general position on quadrics up to automorphisms we get the uniqueness of smooth cubic surfaces up to isomorphism.
Now let us describe the structure of the paper. In Section 2 we study subgroups of order $5$ in the automorphism group of $\mathbb{P}^{1}$ over finite fields of characteristic $2$ In Section 3 we study quadric surfaces and orbits of points on quadric surfaces under the action of the group of order $5$ In Section 4 we study elements and subgroups of Weyl group $W(\mathrm{E}_{6}).$ In Section 5 we recall some basic facts about smooth cubic surfaces and discuss their automorphisms. In Section 6 we prove Theorems 1.4 and 1.5. In Section 7 we discuss the Fermat cubic surface over finite fields of characteristic $2$ and prove Proposition 1.6 and Corollary 1.7.
Notation. Let $X$ be a variety defined over a field $\mathbf{F}.$ If $\mathbf{F}\subset\mathbf{L}$ is an extension of $\mathbf{F},$ then we will denote by $X_{\mathbf{L}}$ the variety
$X_{\mathbf{L}}=X\times_{\mathrm{Spec}(\mathbf{F})}\mathrm{Spec}(\mathbf{L})$
over $\mathbf{L}.$ By $\overline{\mathbf{F}}$ we denote the algebraic closure of $\mathbf{F}.$ By $\mathbf{F}^{sep}$ we denote the separable closure of $\mathbf{F}.$ By $X(\mathbf{F})$ we denote $\mathbf{F}$-points on $X.$ By $\mathfrak{S}_{n}$ we denote the symmetric group of degree $n,$ and by $\mathfrak{A}_{n}$ we denote the alternating group of degree $n.$
Acknowledgment. It is a pleasure for the author to express gratitude to C. A. Shramov for his care, for proposing this problem, a lot of discussion and constant attention. Also the author would like to thank A. S. Trepalin who taught the author his vision of cubic surfaces and their symmetries.
This work was performed at the Steklov International Mathematical Center and supported by the Ministry of Science and Higher Education of the Russian Federation (agreement no. 075-15-2022-265). The author is a winner of the all-Russia mathematical August Moebius contest of graduate and undergraduate student papers and thanks the jury and the board for the high praise of his work. The author was partially supported by Theoretical Physics and Mathematics Advancement Foundation “BASIS” and by the HSE University Basic Research Program.
2. Projective line
In this section we prove some auxiliary lemmas concerning the automorphism group of the projective line over a finite field of characteristic $2$ First of all, let us add the following obvious remark.
Remark 2.1.
The field $\mathbb{F}_{2^{k}}$ contains a non-trivial fifth root of unity if and only if $k=4l.$ Indeed, the former property is equivalent to the condition
$2^{k}-1\equiv 0\mod 5$
This holds if and only if $k=4l.$
Let us give another obvious lemma about the fifth root of unity in finite fields of characteristic $2$
Lemma 2.2.
Let $\xi\in\overline{\mathbb{F}}_{2}$ be a non-trivial fifth root of unity. Then
  1. (i)
    the elements $\xi^{2}+\xi^{3}$ and $\xi+\xi^{4}$ lie in $\mathbb{F}_{4^{k}};$
  2. (ii)
    the elements $\xi^{2}+\xi^{3}$ and $\xi+\xi^{4}$ do not lie in $\mathbb{F}_{2^{2k+1}};$
  3. (iii)
    if $\psi\in\mathrm{Gal}(\mathbb{F}_{4^{2k+1}}/\mathbb{F}_{2^{2k+1}})\simeq\mathbb{Z}/2\mathbb{Z}$ is the generator of the Galois group, then
    $\psi(\xi+\xi^{4})=\xi^{2}+\xi^{3}.$
Proof.
First of all note that $4^{k}\equiv(-1)^{k}\mod 5$ Using this we can prove assertions (i) and (ii). Indeed, assertion (i) follows from the equations
$(\xi+\xi^{4})^{4^{k}}=\xi^{4^{k}}+\xi^{4^{k+1}}=\xi^{(-1)^{k}}+\xi^{(-1)^{k+1}}=\xi+\xi^{4}$
$(\xi^{2}+\xi^{3})^{4^{k}}=\xi^{2\cdot 4^{k}}+\xi^{3\cdot 4^{k}}=\xi^{2\cdot(-1)^{k}}+\xi^{3\cdot(-1)^{k}}=\xi^{2}+\xi^{3}.$
Similarly, assertion (ii) follows from the equations
$(\xi+\xi^{4})^{2^{2k+1}}=\xi^{2\cdot 4^{k}}+\xi^{8\cdot 4^{k}}=\xi^{2\cdot(-1)^{k}}+\xi^{8\cdot(-1)^{k+1}}=\xi^{2}+\xi^{3}$
$(\xi^{2}+\xi^{3})^{2^{2k+1}}=\xi^{4\cdot 4^{k}}+\xi^{6\cdot 4^{k}}=\xi^{(-1)^{k+1}}+\xi^{(-1)^{k}}=\xi+\xi^{4}.$
By definition, $\psi(x)=x^{2^{2k+1}}$ for $x\in\mathbb{F}_{4^{2k+1}}.$ Therefore, assertion (iii) immediately follows from assertions (i) and (ii).
We shall now begin studying the automorphism group of projective line. We give some lemmas about subgroups of order $5$ in $\mathrm{PGL}_{2}(\mathbf{F})$ for a finite field $\mathbf{F}.$ First of all, let us start with the following obvious lemma.
Lemma 2.3.
The order of the group $\mathrm{PGL}_{2}(\mathbb{F}_{2^{2k+1}})$ is not divisible by $5$
Proof.
The order of this group $\mathrm{PGL}_{2}(\mathbb{F}_{2^{2k+1}})$ is equal to
$2^{2k+1}(4^{2k+1}-1)\equiv(-1)^{k}\mod 5$
Lemma 2.4.
Let $\mathbf{F}$ be a finite field and consider the exact sequence
[…]
If $a\in\mathrm{PGL}_{2}(\mathbf{F})$ satisfies $a^{5}=e,$ where $e$ is the neutral element in $\mathrm{PGL}_{2}(\mathbf{F}),$ then there is a matrix $M\in\rho^{-1}(a)$ such that $M^{5}=E,$ where $E$ is the identity matrix.
Proof.
Let us consider some matrix $M\in\rho^{-1}(a).$ We have
(2.1) $M^{5}=c\cdot E,$
where $c\in\mathbf{F}^{*}.$
We divide the proof into two cases. First of all, consider the case when $\mathbf{F}$ does not contain non-trivial fifth root of unity. Then the homomorphism
$f\colon\mathbf{F}^{*}\to\mathbf{F}^{*}$
which sends the element $x\in\mathbf{F}^{*}$ to $f(x)=x^{5}$ is an isomorphism due to triviality of the kernel of $f$ and finiteness of $\mathbf{F}^{*}.$ So we can multiply $M$ by some scalar $m\in\mathbf{F}^{*}$ such that
$m^{5}M^{5}=E.$
Now consider the case when $\mathbf{F}$ contains a non-trivial fifth root of unity. In this case $f$ is not an isomorphism. This means that in equality (2.1) the constant $c$ is not necessary contained in the image of $f.$ If $c$ lies in the image of $f,$ like in the first case we can find $m\in\mathbf{F}^{*}$ such that $m^{5}M^{5}=E$ and we are done. So assume that $c$ does not lie in the image of $f.$ The minimal polynomial $g(x)\in\mathbf{F}[x]$ of $M$ has to be of degree $2,$ because $5]{c}\xi\notin\mathbf{F}^{*}$ for any fifth root of unity $\xi\in\mathbf{F}.$ Besides, $g(x)$ divides $x^{5}-c.$ Therefore, it has to be of the form
[…]
where $i,j\in\{0,1,2,3,4\}$ and $i\neq j.$ However, $5]{c^{2}}\xi^{i+j}$ does not lie in $\mathbf{F}.$ So in this case the matrix $M\in\mathrm{GL}_{2}(\mathbf{F})$ which satisfies the equality (2.1) does not exist. This means that $c$ should lie in the image of $f.$ So as in the previous case we get that $\widetilde{M}^{5}=E$ for some $\widetilde{M}\in\rho^{-1}(a).$
Corollary 2.5.
Let $\mathbf{F}$ be a finite field and consider the exact sequence
[…]
Then for any subgroup $G\simeq\mathbb{Z}/5\mathbb{Z}$ in $\mathrm{PGL}_{2}(\mathbf{F})$ there is a subgroup $\widetilde{G}\simeq\mathbb{Z}/5\mathbb{Z}$ in $\mathrm{GL}_{2}(\mathbf{F})$ such that $\rho(\widetilde{G})=G.$ In addition, if $\mathbf{F}$ contains a non-trivial fifth root of unity $\xi,$ then up to conjugation the group $\widetilde{G}$ can be generated by $\bigl{(}\begin{smallmatrix}1&0\\ 0&\xi\end{smallmatrix}\bigr{)}.$ In particular, the subgroup $G$ is unique up to conjugation.
Proof.
This immediately follows from the fact that by Lemma 2.4 for the group
$G\subset\mathrm{PGL}_{2}(\mathbf{F}),$
which is isomorphic to $\mathbb{Z}/5\mathbb{Z},$ there is a matrix $M\in\mathrm{GL}_{2}(\mathbf{F})$ of order $5$ such that $\rho(M)$ is a generator of $G.$ Since $\xi$ lies in $\mathbf{F}$ the matrix $M$ is diagonalizable. So up to conjugation and multiplication by $\xi^{r}$ for an integer $r$ we have $M=\bigl{(}\begin{smallmatrix}1&0\\ 0&\xi\end{smallmatrix}\bigr{)}.$ This means that the group of order $5$ in $\mathrm{PGL}_{2}(\mathbf{F})$ is unique up to conjugation.
Lemma 2.6.
Assume that $\mathbb{F}_{4^{k}}$ does not contain non-trivial fifth root of unity. Then a subgroup $G$ of order $5$ in the group $\mathrm{PGL}_{2}(\mathbb{F}_{4^{k}})$ is generated up to conjugation by the matrix
(2.2) $\begin{pmatrix}\xi+\xi^{4}&1\\ 1&0\end{pmatrix},$
where $\xi\in\overline{\mathbb{F}}_{2}$ is a non-trivial fifth root of unity. In particular, the subgroup $G$ is unique up to conjugation.
Proof.
Let $\sigma$ be a generator of the group $G\subset\mathrm{PGL}_{2}(\mathbb{F}_{4^{k}})$ of order $5$ By Corollary 2.5 there is a matrix $M\in\mathrm{GL}_{2}(\mathbb{F}_{4^{k}})$ of order $5$ such that its image in $\mathrm{PGL}_{2}(\mathbb{F}_{4^{k}})$ by the natural projection is $\sigma.$ So the matrix $M$ should satisfy the equation
(2.3) $x^{5}-1=0.$
As the minimal polynomial of $M$ divides (2.3) and $M$ is not a scalar matrix, the degree of this minimal polynomial is equal to $2$ Let us fix a non-trivial fifth root of unity $\xi.$ We have
$x^{5}-1=(x-1)(x^{2}+(\xi+\xi^{4})x+1)(x^{2}+(\xi^{2}+\xi^{3})x+1).$
By Lemma 2.2(i) both $\xi+\xi^{4}$ and $\xi^{2}+\xi^{3}$ lie in $\mathbb{F}_{4^{k}}.$ Therefore, the matrix $M$ is conjugate to one of the matrices
$\begin{pmatrix}\xi+\xi^{4}&1\\ 1&0\end{pmatrix}\;\text{or}\;\begin{pmatrix}\xi^{2}+\xi^{3}&1\\ 1&0\end{pmatrix},$
where the minimal polynomial of the first matrix is $x^{2}+(\xi+\xi^{4})x+1$ and the minimal polynomial of the second matrix is $x^{2}+(\xi^{2}+\xi^{3})x+1.$ Note that if $M$ is conjugate to $\bigl{(}\begin{smallmatrix}\xi+\xi^{4}&1\\ 1&0\end{smallmatrix}\bigl{)},$ then $M^{2}$ is conjugate to $\bigl{(}\begin{smallmatrix}\xi^{2}+\xi^{3}&1\\ 1&0\end{smallmatrix}\bigl{)}$ and vice versa. Thus, up to conjugation the group of order $5$ in $\mathrm{PGL}_{2}(\mathbb{F}_{4^{k}})$ can be generated by the matrix (2.2), and so up to conjugation such group is unique.
3. Quadric surfaces
In this section we study smooth quadric surfaces in $\mathbb{P}^{3}$ over finite fields of characteristic $2$ First of all, we refer the reader to [7, §8.1] or [14, Section 1.3] for the definition of Weil restriction of scalars. We recall the result from the paper [11] about the classification of quadric surfaces in $\mathbb{P}^{3}.$
Lemma 3.1 (cf. [11, Lemma 7.3]).
Let $\mathbf{F}$ be an arbitrary field. Let $X$ be a del Pezzo surface of degree $K_{X}^{2}=8$ such that $X_{\overline{\mathbf{F}}}\simeq\mathbb{P}^{1}_{\overline{\mathbf{F}}}\times\mathbb{P}^{1}_{\overline{\mathbf{F}}}.$ Then the surface $X$ is isomorphic either to $C\times C^{\prime},$ where $C$ and $C^{\prime}$ are two conics over $\mathbf{F},$ or to $\mathrm{R}_{\mathbf{L}/\mathbf{F}}(D),$ where $\mathbf{F}\subset\mathbf{L}$ is a separable quadratic extension of $\mathbf{F}$ and $D$ is a conic over $\mathbf{L}.$
The surface $X$ is isomorphic to a quadric surface in $\mathbb{P}^{3}$ if and only if $X\simeq C\times C$ or to $\mathrm{R}_{\mathbf{L}/\mathbf{F}}(C_{\mathbf{L}}),$ for some conic $C$ over $\mathbf{F}.$ In this case one has
$\mathrm{Aut}(X)\simeq\mathrm{Aut}^{\circ}(X)\rtimes\mathbb{Z}/2\mathbb{Z},$
[…]
$\mathrm{Aut}^{\circ}(C\times C)\simeq\mathrm{Aut}(C)\times\mathrm{Aut}(C)$
and
$\mathrm{Aut}^{\circ}(\mathrm{R}_{\mathbf{L}/\mathbf{F}}(C_{\mathbf{L}}))\simeq\mathrm{Aut}(C_{\mathbf{L}}).$
However, as we are dealing with finite fields, by Chevalley–Warning theorem all conics are isomorphic to $\mathbb{P}^{1}$. So we modify Lemma 3.1 for the case when the field is finite.
Lemma 3.2.
Let $X$ be a del Pezzo surface of degree $8$ over a finite field $\mathbf{F}$ such that $X_{\overline{\mathbf{F}}}\simeq\mathbb{P}^{1}_{\overline{\mathbf{F}}}\times\mathbb{P}^{1}_{\overline{\mathbf{F}}}.$ Then $X$ is a quadric surface.
Moreover, if the Picard number of $X$ is equal to $2,$ then $X\simeq\mathbb{P}^{1}\times\mathbb{P}^{1}$ and
$\mathrm{Aut}(X)\simeq\left(\mathrm{PGL}_{2}(\mathbf{F})\times\mathrm{PGL}_{2}(\mathbf{F})\right)\rtimes\mathbb{Z}/2\mathbb{Z},$
where $\mathbb{Z}/2\mathbb{Z}$ interchanges the elements in the pairs $(a,b)\in\mathrm{PGL}_{2}(\mathbf{F})\times\mathrm{PGL}_{2}(\mathbf{F}).$
Otherwise, if the Picard number of $X$ is equal to $1,$ then $X$ is isomorphic to the Weil restriction of scalars $\mathrm{R}_{\mathbf{L}/\mathbf{F}}(\mathbb{P}^{1}),$ where $\mathbf{F}\subset\mathbf{L}$ is a quadratic extension of $\mathbf{F}.$ The automorphism group of $X$ is
(3.1) $\mathrm{Aut}(X)=\mathrm{PGL}_{2}(\mathbf{L})\rtimes\mathbb{Z}/2\mathbb{Z},$
where $\mathbb{Z}/2\mathbb{Z}$ acts on $\mathrm{PGL}_{2}(\mathbf{L})$ as the Galois group $\mathrm{Gal}\left(\mathbf{L}/\mathbf{F}\right)\simeq\mathbb{Z}/2\mathbb{Z}.$
Remark 3.3.
Let $Q$ be the Weil restriction of scalars $\mathrm{R}_{\mathbf{L}/\mathbf{F}}(\mathbb{P}^{1}),$ where $\mathbf{F}\subset\mathbf{L}$ is a separable quadratic extension of $\mathbf{F}.$ Let $\tau$ be a generator of the Galois group of this extension. Let
$\phi=(\chi,\upsilon)\in\mathrm{PGL}_{2}(\mathbf{L})\rtimes\mathbb{Z}/2\mathbb{Z}$
be an automorphism of $Q.$ Then this automorphism extends to the automorphism
$\phi=(\chi,\tau(\chi),\upsilon)\in\left(\mathrm{PGL}_{2}(\mathbf{L})\times\mathrm{PGL}_{2}(\mathbf{L})\right)\rtimes\mathbb{Z}/2\mathbb{Z}$
of $Q_{\mathbf{L}}=\mathbb{P}^{1}_{\mathbf{L}}\times\tau_{*}\mathbb{P}^{1}_{\mathbf{L}},$ where $\tau_{*}\mathbb{P}^{1}_{\mathbf{L}}\simeq\mathbb{P}^{1}$ is Galois conjugate to $\mathbb{P}^{1}_{\mathbf{L}}.$ This immediately follows from the equivalence of the category of the projective varieties over $\mathbf{F}$ and the category of the projective varieties over $\mathbf{L}$ with descent data.
We need the following definition.
Definition 3.4.
We say that $5$ points $P_{1},\ldots,P_{5}\in\mathbb{P}^{1}\times\mathbb{P}^{1}$ are in general position if no $2$ points of this set lie on a divisor of bidegree $(1,0)$ and $(0,1)$ and no $4$ points of this set lie on a divisor of bidegree $(1,1).$ If $Q\subset\mathbb{P}^{3}_{\mathbf{F}}$ is a quadric surface over a field $\mathbf{F},$ then we say that $5$ points $P_{1},\ldots,P_{5}\in Q$ are in general position if they are in general position on
$Q_{\overline{\mathbf{F}}}\simeq\mathbb{P}^{1}_{\overline{\mathbf{F}}}\times\mathbb{P}^{1}_{\overline{\mathbf{F}}}.$
Now let us study the sets of $5$ points in general position on a quadric surface over $\mathbb{F}_{4^{k}}.$
Proposition 3.5.
Let $P_{1},\ldots,P_{5}$ be a set of points in general position on $\mathbb{P}^{1}\times\mathbb{P}^{1}$ over the field $\mathbb{F}_{4^{2k}}$ such that they form an orbit of some subgroup
$\mathbb{Z}/5\mathbb{Z}\subset\mathrm{Aut}(\mathbb{P}^{1}\times\mathbb{P}^{1}).$
Then the set of points $P_{1},\ldots,P_{5}$ is unique up to automorphisms of $\mathbb{P}^{1}\times\mathbb{P}^{1}.$
Proof.
By Remark 2.1 the field $\mathbb{F}_{4^{2k}}$ contains a non-trivial fifth root of unity. We need to define the embedding
$\eta:\mathbb{Z}/5\mathbb{Z}\hookrightarrow\mathrm{Aut}(\mathbb{P}^{1}\times\mathbb{P}^{1})\simeq\left(\mathrm{PGL}_{2}(\mathbb{F}_{4^{2k}})\times\mathrm{PGL}_{2}(\mathbb{F}_{4^{2k}})\right)\rtimes\mathbb{Z}/2\mathbb{Z}.$
By Corollary 2.5 the projection of the image of the embedding on any copy of the group $\mathrm{PGL}_{2}(\mathbb{F}_{4^{2k}})$ is conjugate to the element $\bigl{(}\begin{smallmatrix}1&0\\ 0&\xi\end{smallmatrix}\bigr{)}$ for a non-trivial fifth root of unity $\xi\in\mathbb{F}_{4^{2k}}.$
By the condition of general position of $P_{1},\ldots,P_{5}$ (i.e. no $2$ points lie on a divisor of bidegree $(1,0)$ and $(0,1)$) we get that $\mathbb{Z}/5\mathbb{Z}$ projects non-trivially in both copies of $\mathrm{PGL}_{2}(\mathbb{F}_{4^{2k}}).$ In other words, up to conjugation we have
$\eta(\sigma)=\left\langle\begin{pmatrix}1&0\\ 0&\xi_{1}\end{pmatrix},\begin{pmatrix}1&0\\ 0&\xi_{2}\end{pmatrix},\mathrm{id}\right\rangle,$
where $\sigma$ is a generator of $\mathbb{Z}/5\mathbb{Z}$ and $\xi_{1}$ and $\xi_{2}$ are non-trivial fifth root of unity.
Without loss of generality we can assume that $P_{1}$ is $([1:1],[1:1])\in\mathbb{P}^{1}\times\mathbb{P}^{1}.$ Indeed, any point of the form
(3.2) $([a:b],[c:d])\in\mathbb{P}^{1}\times\mathbb{P}^{1},$
where $a$, $b$, $c$ and $d$ are non-zero, can be obtained from the point ([1:1],[1:1])([1:1],[1:1])( [ 1 : 1 ] , [ 1 : 1 ] ) by the action of the automorphism
$\left\langle\begin{pmatrix}a&0\\ 0&b\end{pmatrix},\begin{pmatrix}c&0\\ 0&d\end{pmatrix},\mathrm{id}\right\rangle.$
This automorphism commutes with $\eta(\sigma).$ If one of $a,$ $b,$ $c,$ $d$ is zero, then the orbit of the point (3.2) is not in general position. By general position of the points $P_{1},$ $P_{2},$ $P_{3},$ $P_{4},$ $P_{5}$ (i.e. no $4$ points lie on a divisor of bidegree $(1,1)$) we get that $\xi_{1}\neq\xi_{2}.$ Let us fix some non-trivial fifth root of unity $\xi.$ So up to conjugation we have the following embeddings:
$\eta_{2}(\sigma)=\left\langle\begin{pmatrix}1&0\\ 0&\xi\end{pmatrix},\begin{pmatrix}1&0\\ 0&\xi^{2}\end{pmatrix},\mathrm{id}\right\rangle,$
$\eta_{3}(\sigma)=\left\langle\begin{pmatrix}1&0\\ 0&\xi\end{pmatrix},\begin{pmatrix}1&0\\ 0&\xi^{3}\end{pmatrix},\mathrm{id}\right\rangle,$
$\eta_{4}(\sigma)=\left\langle\begin{pmatrix}1&0\\ 0&\xi\end{pmatrix},\begin{pmatrix}1&0\\ 0&\xi^{4}\end{pmatrix},\mathrm{id}\right\rangle.$
The orbit of the point $P_{1}$ under the action of $\eta_{2}(\mathbb{Z}/5\mathbb{Z})$ is
(3.3) $([1:1],[1:1]);\;([1:\xi],[1:\xi^{2}]);\;([1:\xi^{2}],[1:\xi^{4}]);\;([1:\xi^{3}],[1:\xi]);\;([1:\xi^{4}],[1:\xi^{3}]).$
The orbit of the point $P_{1}$ under the action of $\eta_{3}(\mathbb{Z}/5\mathbb{Z})$ is
(3.4) $([1:1],[1:1]);\;([1:\xi],[1:\xi^{3}]);\;([1:\xi^{2}],[1:\xi]);\;([1:\xi^{3}],[1:\xi^{4}]);\;([1:\xi^{4}],[1:\xi^{2}]).$
The orbit of the point $P_{1}$ under the action of $\eta_{4}(\mathbb{Z}/5\mathbb{Z})$ is
(3.5) $([1:1],[1:1]);\;([1:\xi],[1:\xi^{4}]);\;([1:\xi^{2}],[1:\xi^{3}]);\;([1:\xi^{3}],[1:\xi^{2}]);\;([1:\xi^{4}],[1:\xi]).$
One can see that the orbit (3.4) can be obtained by the action of $\mathbb{Z}/2\mathbb{Z}$ on (3.3) by interchanging two copies of $\mathbb{P}^{1}.$ Also note that the orbit (3.5) can be obtained from $5$ points
(3.6) $([1:1],[1:1]);\;([1:\xi],[1:\xi]);\;([1:\xi^{2}],[1:\xi^{2}]);\;([1:\xi^{3}],[1:\xi^{3}]);\;([1:\xi^{4}],[1:\xi^{4}])$
by acting with the element
$\left\langle\begin{pmatrix}1&0\\ 0&1\end{pmatrix},\begin{pmatrix}0&1\\ 1&0\end{pmatrix},\mathrm{id}\right\rangle\in\left(\mathrm{PGL}_{2}(\mathbb{F}_{4^{2k}})\times\mathrm{PGL}_{2}(\mathbb{F}_{4^{2k}})\right)\rtimes\mathbb{Z}/2\mathbb{Z}.$
The orbit (3.6) is prohibited, as all points in the orbit lie on a divisor of bidegree $(1,1).$ Thus the orbit (3.5) is also prohibited. Therefore, we get the only allowed orbit in this case.
Proposition 3.6.
Let $P_{1},\ldots,P_{5}$ be a set of different points in general position on $\mathbb{P}^{1}\times\mathbb{P}^{1}$ over the field $\mathbb{F}_{4^{2k+1}}$ such that they form an orbit of some subgroup
$\mathbb{Z}/5\mathbb{Z}\subset\mathrm{Aut}(\mathbb{P}^{1}\times\mathbb{P}^{1}).$
Then the set of points $P_{1},\ldots,P_{5}$ is unique up to automorphisms of $\mathbb{P}^{1}\times\mathbb{P}^{1}.$
Proof.
By Remark 2.1 the field $\mathbb{F}_{4^{2k+1}}$ does not contain non-trivial fifth root of unity. Let $\sigma$ be a generator of $\mathbb{Z}/5\mathbb{Z}.$ By Lemma 2.6 the projection of the image of $\sigma$ under the embedding in $\mathrm{Aut}(\mathbb{P}^{1}\times\mathbb{P}^{1})$ on any copy of $\mathrm{PGL}_{2}(\mathbb{F}_{4^{2k+1}})$ is conjugate to the matrix (2.2). By the condition of general position of $P_{1},\ldots,P_{5}$ (i.e. no $2$ points lie on a divisor of bidegree $(1,0)$ and $(0,1)$) we get that the image of $\sigma$ projects non-trivially in both copies of $\mathrm{PGL}_{2}(\mathbb{F}_{4^{2k+1}}).$
As in the proof of Proposition 3.5 we can assume that
$P_{1}=([1:1],[1:1])\in\mathbb{P}^{1}\times\mathbb{P}^{1}.$
By general position of $P_{1},$ $P_{2},$ $P_{3},$ $P_{4},$ $P_{5}$ (i.e. no $4$ points lie on a divisor of bidegree $(1,1)$) the projections in both copies of $\mathrm{PGL}_{2}(\mathbb{F}_{4^{2k+1}})$ are different. So up to conjugation the image of $\mathbb{Z}/5\mathbb{Z}$ is generated by the following elements in the group $\mathrm{Aut}(\mathbb{P}^{1}\times\mathbb{P}^{1}):$
$\eta(\sigma)=\left\langle\begin{pmatrix}\xi+\xi^{4}&1\\ 1&0\end{pmatrix},\begin{pmatrix}\xi^{2}+\xi^{3}&1\\ 1&0\end{pmatrix},\mathrm{id}\right\rangle,$
$\widetilde{\eta}(\sigma)=\left\langle\begin{pmatrix}\xi^{2}+\xi^{3}&1\\ 1&0\end{pmatrix},\begin{pmatrix}\xi+\xi^{4}&1\\ 1&0\end{pmatrix},\mathrm{id}\right\rangle,$
where $\xi\in\mathbb{F}_{16^{2k+1}}$ is a non-trivial fifth root of unity.
The orbit of $P_{1}$ under the action of $\eta(\sigma)$ can be obtained from the orbit of $P_{1}$ under the action of $\widetilde{\eta}(\sigma)$ by the action of $\mathbb{Z}/2\mathbb{Z}$ which interchanges two copies of $\mathbb{P}^{1}.$ So we get that there is only one allowed orbit up to automorphisms of $\mathbb{P}^{1}\times\mathbb{P}^{1}$ in this case.
Now we study the Weil restriction of scalars of $\mathbb{P}^{1}.$
Proposition 3.7.
Let $Q$ be a quadric surface over the field $\mathbb{F}_{2^{2k+1}}$ which is isomorphic to the Weil restriction of scalars $\mathrm{R}_{\mathbb{F}_{4^{2k+1}}/\mathbb{F}_{2^{2k+1}}}(\mathbb{P}^{1}).$ Let $P_{1},\ldots,P_{5}$ be a set of different points in general position on $Q$ over the field $\mathbb{F}_{2^{2k+1}}$ such that they form an orbit of some subgroup $\mathbb{Z}/5\mathbb{Z}\subset\mathrm{Aut}(Q).$ Then the set of points $P_{1},\ldots,P_{5}$ is unique up to automorphisms of $Q.$
Proof.
Let us study the possible sets of $5$ points in general position on $Q.$ By Lemma 2.6 a generator of $\mathbb{Z}/5\mathbb{Z}$ in $\mathrm{PGL}_{2}(\mathbb{F}_{4^{2k+1}})$ is conjugate to the matrix (2.2) for some non-trivial fifth root of unity $\xi\in\overline{\mathbb{F}}_{2}$. Therefore, we get that up to conjugation the subgroup $\mathbb{Z}/5\mathbb{Z}$ in $\mathrm{Aut}(Q)$ is generated by
$\alpha(\sigma)=\left\langle\begin{pmatrix}\xi+\xi^{4}&1\\ 1&0\end{pmatrix},\mathrm{id}\right\rangle.$
By Remark 3.3 the automorphism $\alpha(\sigma)$ of $Q$ extends to the automorphism
$\alpha(\sigma)_{\mathbb{F}_{4^{2k+1}}}=\left\langle\begin{pmatrix}\xi+\xi^{4}&1\\ 1&0\end{pmatrix},\begin{pmatrix}\xi^{2}+\xi^{3}&1\\ 1&0\end{pmatrix},\mathrm{id}\right\rangle$
on $Q_{\mathbb{F}_{4^{2k+1}}}$ since by Lemma 2.2(iii) the elements $\xi+\xi^{4}$ and $\xi^{2}+\xi^{3}$ in $\mathbb{F}_{4^{2k+1}}$ are Galois conjugate.
Without loss of generality we can assume that $P_{1}$ is $([1:1],[1:1])\in Q_{\mathbb{F}_{2^{2k+1}}}.$ Indeed, any point of the form
(3.7) $([a:b],[\psi(a):\psi(b)])\in\mathbb{P}^{1}\times\mathbb{P}^{1},$
where $a$ and $b$ are non-zero and $\psi\in\mathrm{Gal}(\mathbb{F}_{4^{2k+1}}/\mathbb{F}_{2^{2k+1}})$, can be obtained from the point ([1:1],[1:1])([1:1],[1:1])( [ 1 : 1 ] , [ 1 : 1 ] ) by the action of the automorphism
$\left\langle\begin{pmatrix}a&0\\ 0&b\end{pmatrix},\mathrm{id}\right\rangle.$
This automorphism commutes with $\alpha(\sigma).$ If $a$ or $b$ is zero, then the orbit of the point (3.7) is not in general position.
Let us consider the orbit of the point
$P_{1}=([1:1],[1:1])\in Q_{\mathbb{F}_{2^{2k+1}}}$
under the action of the group generated by $\alpha(\sigma)_{\mathbb{F}_{4^{2k+1}}}.$ It is
$([1:1],[1:1]);\;([\xi^{2}+\xi^{3}:1],[\xi+\xi^{4}:1]);\;([0:1],[0:1]);$
$([1:0],[1:0]);\;([\xi+\xi^{4}:1],[\xi^{2}+\xi^{3}:1]).$
It is not hard to see that the points of this quintuple are in general position. Therefore, we get that there is only one orbit of $5$ points in general position on $Q$ up to the action of the group $\mathrm{Aut}(Q)\simeq\mathrm{PGL}_{2}(\mathbb{F}_{4^{2k+1}})\rtimes\mathbb{Z}/2\mathbb{Z}.$
4. The Weyl group of type $\mathrm{E}_{6}$
In this section we prove some lemmas concerning the Weyl group $W(\mathrm{E}_{6}).$ At the beginning we recall the structure of the group $W(\mathrm{E}_{6}).$
Lemma 4.1 ([15, §3.12.4]).
For the Weyl group $W(\mathrm{E}_{6})$ we have
$W(\mathrm{E}_{6})\simeq\mathrm{PSU}_{4}(\mathbb{F}_{2})\rtimes\mathbb{Z}/2\mathbb{Z}$
with a non-trivial action of $\mathbb{Z}/2\mathbb{Z}$ on $\mathrm{PSU}_{4}(\mathbb{F}_{2}).$
Lemma 4.2.
The subgroup $\mathrm{PSU}_{4}(\mathbb{F}_{2})$ in $W(\mathrm{E}_{6})$ coincides with the subgroup of elements with determinant equal to $1$ considered as the action on the root system $\mathrm{E}_{6}.$
Proof.
As all elements in $W(\mathrm{E}_{6})$ are isometries their determinant is ±$\pm 1.$ So the subgroup of elements with determinant equal to $1$ is a normal subgroup of index $2$ since it is a kernel of the surjection $W(\mathrm{E}_{6})\twoheadrightarrow\mathbb{Z}/2\mathbb{Z}.$ Let $K$ be the kernel of this homomorphism. Then $H=\mathrm{PSU}_{4}(\mathbb{F}_{2})\cap K$ is a normal subgroup in $\mathrm{PSU}_{4}(\mathbb{F}_{2}).$ By [15, §3.6.1] the group $\mathrm{PSU}_{4}(\mathbb{F}_{2})$ is simple, thus $H$ is either trivial, or coincides with $\mathrm{PSU}_{4}(\mathbb{F}_{2}).$ However, $H$ is not trivial since $g^{2}\in\mathrm{PSU}_{4}(\mathbb{F}_{2})$ for any element $g\in K.$ Therefore, $\mathrm{PSU}_{4}(\mathbb{F}_{2})$ coincides with $K.$
Now we recall the results from [1, Table 9] and [12, Theorem 1] about the conjugacy classes of the elements of order $2$, $5$ and $10$ in the Weyl group $W(\mathrm{E}_{6}).$
Lemma 4.3.
In Table 1 we list conjugacy classes of elements in the group $W(\mathrm{E}_{6})$ of orders $2$, $5$ and $10,$ their cardinalities, orders of their centralizers, collections of eigenvalues of the representation on the root system $\mathrm{E}_{6}.$
Order of Conjugacy class Cardinality Order of Eigenvalues
element centralizer
$2$ $A_{1}$ $36$ $4401$$,1,1,1,1,1$ 1
$2$ $A_{1}^{2}$ $270$ $921$$,-1,1,1,1,1$ 1
$2$ $A_{1}^{3}$ $540$ $96$ $-1,-1,-1,1,1,1$
$2$ $A_{1}^{4}$ $45$ $1521$$,-1,-1,-1,1,1$ 1
$5$ $A_{4}$ $5184$ $10$ $\xi,\xi^{2},\xi^{3},\xi^{4},1,1$
$10$ $A_{4}\times A_{1}$ $5184$ $10$ $\xi,\xi^{2},\xi^{3},\xi^{4},-1,1$
Table 1. Table of conjugacy classes of elements of order $2$, $5$ and $10$ in $W(\mathrm{E}_{6})$.
Here $\xi\in\mathbb{C}$ is a non-trivial fifth root of unity.
Lemma 4.4.
The centralizer of an element of order $5$ in $W(\mathrm{E}_{6})$ is isomorphic to $\mathbb{Z}/10\mathbb{Z}.$
Proof.
As by Lemma 4.3 the order of the centralizer of an element $g$ of order $5$ in $W(\mathrm{E}_{6})$ is equal to $10$ This means that the centralizer of $g$ in the Weyl group is isomorphic either to $\mathbb{Z}/10\mathbb{Z},$ or to a dihedral group of order $10,$ since these two groups are the only groups of order $10$ up to isomorphism. However, the second case is impossible because $g$ lies in its centralizer.
Lemma 4.5.
Let $\tau\in W(\mathrm{E}_{6})$ be an element of order $10$ Then $\tau^{5}$ lies in the conjugacy class $A_{1}.$
Proof.
This directly follows from Lemma 4.3. By Table 1 we see that the eigenvalues of an element lying in $A_{4}\times A_{1}$ are
$\xi,\xi^{2},\xi^{3},\xi^{4},-1,1,$
where $\xi\in\mathbb{C}$ is a non-trivial fifth root of unity. We have $g=\tau^{5}.$ So the eigenvalues of $\tau^{5}$ are
$-1,1,1,1,1,1$
which are the eigenvalues of elements in conjugacy class $A_{1}.$
Corollary 4.6.
Let $g\in W(\mathrm{E}_{6})$ be an element of order $2$ which lies in the centralizer of an element of order $5$ Then $g$ lies in the conjugacy class $A_{1}.$
Proof.
Let $h$ be an element of order $5$ in $W(\mathrm{E}_{6}).$ Then as $g$ is in the centralizer of $h$ the element $gh$ is of order $10$ By Lemma 4.3 there is only one conjugacy class $A_{4}\times A_{1}$ of elements of order $10$ in $W(\mathrm{E}_{6}).$ Therefore, by Lemma 4.5 the element $g$ of order $2$ lies in the conjugacy class $A_{1}.$
Now we study subgroups in $\mathrm{PSU}_{4}(\mathbb{F}_{2})$ and in the Weyl group $W(\mathrm{E}_{6}).$
Lemma 4.7 ([2, p. 26]).
The maximal proper subgroups in $\mathrm{PSU}_{4}(\mathbb{F}_{2})$ are listed in Table 2.
Subgroup Order
$\left(\mathbb{Z}/2\mathbb{Z}\right)^{4}\rtimes\mathfrak{A}_{5}$ $960$
$\mathfrak{S}_{6}$ $720$
[…]
$\left(\mathbb{Z}/3\mathbb{Z}\right)^{3}\rtimes\mathfrak{S}_{4}$ $648$
[…]
Table 2. Table of maximal subgroups in $\mathrm{PSU}_{4}(\mathbb{F}_{2})$.
[…]
Lemma 4.8.
The subgroup in $W(\mathrm{E}_{6})$ isomorphic to $\mathfrak{A}_{6}$ is unique up to conjugation.
Proof.
First of all, let us prove that the subgroup of $W(\mathrm{E}_{6})$ isomorphic to $\mathfrak{A}_{6}$ lies in $\mathrm{PSU}_{4}(\mathbb{F}_{2}).$ Let us consider the canonical projection
$p:W(\mathrm{E}_{6})\simeq\mathrm{PSU}_{4}(\mathbb{F}_{2})\rtimes\mathbb{Z}/2\mathbb{Z}\to\mathbb{Z}/2\mathbb{Z}.$
Then the kernel of the homomorphism
$p\mid_{\mathfrak{A}_{6}}:\mathfrak{A}_{6}\to\mathbb{Z}/2\mathbb{Z}$
is the whole group $\mathfrak{A}_{6}$ since it is simple. This means that this $\mathfrak{A}_{6}$ lies in $\mathrm{PSU}_{4}(\mathbb{F}_{2}).$
Now let us prove that in $\mathrm{PSU}_{4}(\mathbb{F}_{2})$ there is the only subgroup isomorphic to $\mathfrak{A}_{6}$ up to conjugation. The subgroup in $\mathrm{PSU}_{4}(\mathbb{F}_{2})$ isomorphic to $\mathfrak{A}_{6}$ should lie in some maximal subgroup of $\mathrm{PSU}_{4}(\mathbb{F}_{2}).$ By Lemma 4.7 the only maximal subgroup, the order of which is divisible by the order of $\mathfrak{A}_{6},$ is isomorphic to $\mathfrak{S}_{6}.$ So $\mathfrak{A}_{6}$ has to lie in $\mathfrak{S}_{6}.$ Moreover, the subgroup in $\mathfrak{S}_{6}$ isomorphic to $\mathfrak{A}_{6}$ is unique. Therefore, as $\mathfrak{S}_{6}$ is unique up to conjugation in $\mathrm{PSU}_{4}(\mathbb{F}_{2}),$ we get that $\mathfrak{A}_{6}$ is unique in $\mathrm{PSU}_{4}(\mathbb{F}_{2})$ up to conjugation. As a consequence, there is the only $\mathfrak{A}_{6}$ in $W(\mathrm{E}_{6})$ up to conjugation.
Lemma 4.9.
The centralizers of the subgroups $\mathrm{PSU}_{4}(\mathbb{F}_{2})$ and
$\left(\mathbb{Z}/2\mathbb{Z}\right)^{4}\rtimes\mathfrak{A}_{5}\subset\mathrm{PSU}_{4}(\mathbb{F}_{2})$
in the Weyl group $W(\mathrm{E}_{6})$ are trivial.
Proof.
Let us consider the subgroup $G\simeq\left(\mathbb{Z}/2\mathbb{Z}\right)^{4}\rtimes\mathfrak{A}_{5}$ in $\mathrm{PSU}_{4}(\mathbb{F}_{2}).$ It is enough to prove that the centralizer of $G$ in $W(\mathrm{E}_{6})$ is trivial. There is an element of order $5$ in $G.$ Therefore, by Lemma 4.4 the centraliser of $G$ lies in $\mathbb{Z}/10\mathbb{Z}.$ However, the element of order $5$ does not lie in the center of $G.$ This means that the centralizer of $G$ lies in $\mathbb{Z}/2\mathbb{Z}.$ Assume that the centralizer is isomorphic to $\mathbb{Z}/2\mathbb{Z}$ and let $h$ be its generator. By Lemma 4.4 the element $h$ lies in the conjugacy class $A_{1}.$ By Lemma 4.3 the order of its centralizer is equal to $1440$ But the order $|G|=960$ does not divide $1440$ Therefore, the centralizer of $G$ is trivial. Moreover, from this we get that the centralizer of $\mathrm{PSU}_{4}(\mathbb{F}_{2})$ is also trivial.
5. Cubic surfaces
In this section we introduce the notation related to smooth cubic surfaces and recall the description of their geometry. More details on smooth cubic surfaces can be found in [5, Chapter 9] or [10, Chapter \@slowromancapiv@].
Let $S\subset\mathbb{P}^{3}$ be a smooth cubic surface over a field $\mathbf{F}.$ Then $S_{\overline{\mathbf{F}}}$ is a blowup of $6$ points
$P_{1},\;P_{2},\;P_{3},\;P_{4},\;P_{5},\;P_{6}$
in general position on $\mathbb{P}^{2}_{\overline{\mathbf{F}}}.$ In other words, there is a birational morphism
$\pi\colon S_{\overline{\mathbf{F}}}\to\mathbb{P}^{2}_{\overline{\mathbf{F}}}$
which blows down the preimages of points $P_{1},\ldots,P_{6}.$ Denote by
$E_{1},\;E_{2},\;E_{3},\;E_{4},\;E_{5},\;E_{6}$
the preimages of $P_{1},$ $P_{2},$ $P_{3},$ $P_{4},$ $P_{5}$ and $P_{6},$ respectively. The exceptional curves are lines on $S_{\overline{\mathbf{F}}}\subset\mathbb{P}^{3}_{\overline{\mathbf{F}}}.$ The self-intersection of any line on a smooth cubic surface is equal to $-1.$ It is well known that there are exactly $27$ lines on a smooth cubic surface over an algebraically closed field. We denote the other $21$ lines by $Q_{i}$ and $L_{ij}=L_{ji}$ for $i,j\in\{1,\ldots,6\}$ and $i\neq j.$ Denote by $H$ the pullback of the class of a line on $\mathbb{P}^{2}_{\overline{\mathbf{F}}}.$ Then the $27$ lines on the cubic surface $S_{\overline{\mathbf{F}}}$ are the following effective divisors
$E_{i}\quad\text{for}\quad i\in\{1,\ldots,6\};$
$L_{ij}\sim H-E_{i}-E_{j}\quad\text{for}\quad i\neq j;\;i,j\in\{1,\ldots,6\};$
$Q_{i}\sim 2H+E_{i}-\sum_{j=1}^{6}E_{j}\quad\text{for}\quad i\in\{1,\ldots,6\}.$
Their intersections are:
$E_{i}^{2}=Q_{i}^{2}=L_{ij}^{2}=-1\quad\text{for}\quad i\neq j;$
$E_{i}\cdot E_{j}=Q_{i}\cdot Q_{j}=0\quad\text{for}\quad i\neq j;$
$E_{i}\cdot Q_{j}=L_{ij}\cdot E_{i}=L_{ij}\cdot Q_{i}=1\quad\text{for}\quad i\neq j;$
$L_{ij}\cdot L_{kl}=1,\quad\text{where}\;i,j,k,l\in\{1,\ldots,6\}\;\text{are pairwise distinct;}$;
$L_{ij}\cdot L_{jk}=0,\quad\text{where}\;i,j,k\in\{1,\ldots,6\}\;\text{are pairwise distinct}.$
Let us consider the set
$\Delta=\{\alpha\in\mid K_{S_{\overline{\mathbf{F}}}}\cdot\alpha=0\;\text{and}\;\alpha^{2}=-2\}.$
Then this set is the root system $\mathrm{E}_{6}.$ The Weyl group $W(\mathrm{E}_{6})$ acts on the Picard group $\mathrm{Pic}(S_{\overline{\mathbf{F}}})$ by orthogonal transformations with respect to the intersection form. It fixes $K_{S_{\overline{\mathbf{F}}}}$ and preserves the cone of effective divisors.
Recall the following important theorem.
Theorem 5.1 (see, for instance, [5, Corollary 8.2.40]).
The automorphism group of a smooth cubic surface $S$ is isomorphic to a subgroup in $W(\mathrm{E}_{6}).$
Although in the literature Theorem 5.1 is proved over fields of characteristic zero, exactly the same proof also works in the case of positive characteristic.
Using the classification of automorphism groups of smooth cubic surfaces we can say more about the group, where automorphism groups lie.
Lemma 5.2.
Let $S$ be a smooth cubic surface. Then $\mathrm{Aut}(S)\subseteq\mathrm{PSU}_{4}(\mathbb{F}_{2}).$
Proof.
This directly follows from the classification of the automorphism groups which can be found in [8, Theorem 5.3] in zero characteristic and in [6, Theorem 1.1] for positive characteristic.
Let us go back to the conjugacy classes of elements in $W(\mathrm{E}_{6})$ in the light of the action on the Picard group on a cubic surface.
Lemma 5.3 ([3, Table 7.1, column 8]).
Let $S$ be a smooth cubic surface over an algebraically closed field. Consider an element $g\in W(\mathrm{E}_{6})$ of order $2$ acting on $S.$ The following assertions hold.
  1. (i)
    Assume that $g$ lies in the conjugacy class $A_{1}.$ Then there are $15$ lines on $S$ which are invariant under the action of $g.$ Moreover, there are $6$ pairs of skew lines on $S$ which are invariant under the action of $g$ and in any pair the lines are interchanged by $g.$
  2. (ii)
    Assume that $g$ lies in the conjugacy class $A_{1}^{2}.$ Then there are $7$ lines on $S$ which are invariant under the action of $g.$ There are $8$ pairs of skew lines on $S$ which are invariant under the action of $g$ and in any pair the lines are interchanged by $g.$ Moreover, there are $2$ pairs of intersecting lines on $S$ which are invariant under the action of $g$ and in any pair the lines are interchanged by $g.$
  3. (iii)
    Assume that $g$ lies in the conjugacy class $A_{1}^{3}.$ Then there are $3$ lines which are invariant under the action of $g.$ There are $6$ pairs of skew lines on $S$ which are invariant under the action of $g$ and in any pair the lines are interchanged by $g.$ Moreover, there are $6$ pairs of intersecting lines on $S$ which are invariant under the action of $g$ and in any pair the lines are interchanged by $g.$
  4. (iv)
    Assume that $g$ lies in the conjugacy class $A_{1}^{4}.$ Then there are $3$ lines which are invariant under the action of $g.$ Moreover, there are $12$ pairs of intersecting lines which are invariant under the action of $g$ and in any pair the lines are interchanged by $g.$
Remark 5.4.
Let $S$ be a smooth cubic surface over a field $\mathbf{F}.$ Note that the natural action of the group $G\simeq\mathfrak{S}_{6}$ on the Picard group of $S_{\overline{\mathbf{F}}}$ which permutes the indices of exceptional curves $E_{1},\ldots,E_{6}$ in not geometric, i.e. $G$ does not lie in the automorphism group of $S_{\overline{\mathbf{F}}}.$ Suppose the contrary. Let us consider the transposition $(12).$ This element fixes the following lines
(5.1) $L_{12},\,L_{34},\,L_{35},\,L_{36},\,L_{45},\,L_{46},\,L_{56},\,E_{3},\,E_{4},\,E_{5},\,E_{6},\,Q_{3},\,Q_{4},\,Q_{5},\,Q_{6}.$
therefore, by Lemma 5.3 the transposition lies in conjugacy class $A_{1}.$ By Lemma 4.3 its determinant is equal to $-1.$ So by Lemma 4.2 the transposition does not lie in $\mathrm{PSU}_{4}(\mathbb{F}_{2}).$
Remark 5.5.
Let us consider an element $\sigma$ of order $2$ in $W(\mathrm{E}_{6})$ which interchanges $E_{i}$ and $Q_{i}$ for all $i\in\{1,\ldots,6\}$ and fixes $L_{ij}$ for all $i,j\in\{1,\ldots,6\}$ and $i\neq j.$ Then $\sigma$ lies in the conjugacy class of type $A_{1}$ by Lemma 5.3 since $\sigma$ fixes $15$ lines.
Now let us study the blowdown of a smooth cubic surface to the quadric in the light of automorphism groups. From now on we use the following definition.
Definition 5.6.
Let $S$ be a smooth cubic surface over a field $\mathbf{F}.$ We say that $5$ skew lines $l_{1},\ldots,l_{5}$ on $S$ are in quadratic position if blowing them down over $\overline{\mathbf{F}}$ gives $\mathbb{P}^{1}_{\overline{\mathbf{F}}}\times\mathbb{P}^{1}_{\overline{\mathbf{F}}}.$
We emphasize that in Definition 5.6 it is required that each of the lines $l_{1},\ldots,l_{5}$ is defined over $\mathbf{F}.$
Remark 5.7.
By [9, Chapter 4, Proposition 2.7 and Exercise 2.9] the blowup of points in general position on a quadric surface $\mathbb{P}^{1}_{\overline{\mathbf{F}}}\times\mathbb{P}^{1}_{\overline{\mathbf{F}}}$ gives a smooth cubic surface. In other words, the preimages of points in general position on a quadric surface are the lines in quadratic position. Conversely, the image of the blowdown of lines in quadratic position on a smooth cubic surface is a set of points in general position on $\mathbb{P}^{1}_{\overline{\mathbf{F}}}\times\mathbb{P}^{1}_{\overline{\mathbf{F}}}.$
Lemma 5.8.
Let $S$ be a smooth cubic surface over an algebraically closed field $\mathbf{F}.$ Then the lines
(5.2) $L_{12},\;L_{13},\;L_{14},\;L_{15},\;L_{16}.$
are in quadratic position.
Proof.
By definition, the lines (5.2) are pairwise skew. Let us blow down the first four lines in (5.2). We obtain a morphism $f\colon S\to\widehat{S},$ where $\widehat{S}$ is a del Pezzo surface of degree $7$ and $f$ is the blowdown of $L_{12}$, $L_{13}$, $L_{14}$ and $L_{15}$. There are three $(-1)$-curves on $\widehat{S},$ namely, $f(L_{16}),$ $f(E_{6})$ and $f(Q_{6}).$ Moreover, we have the following intersection numbers:
$f(E_{6})\cdot f(Q_{6})=0\quad\text{and}\quad f(L_{16})\cdot f(E_{6})=f(L_{16})\cdot f(Q_{6})=1.$
So if we blow down $f(L_{16}),$ we get a del Pezzo surface of degree $8$ without $(-1)$-curves. This means that we get $\mathbb{P}^{1}\times\mathbb{P}^{1}.$
Let $S$ be a smooth cubic surface over a field $\mathbf{F}.$ Let us consider a Galois extension $\mathbf{L}$ of the field $\mathbf{F}.$ By [10, Theorem 23.8(ii)] the Galois group $\mathrm{Gal}(\mathbf{L}/\mathbf{F})$ acts on $\mathrm{Pic}(S_{\mathbf{L}})$ and preserves the canonical class of $S$ and the intersection form. This means that there is a homomorphism
$\mathrm{Gal}(\mathbf{L}/\mathbf{F})\to\mathrm{Aut}(\mathrm{Pic}(S_{\mathbf{L}}))\subset W(\mathrm{E}_{6}).$
Let $\Gamma$ be the image of the Galois group $\mathrm{Gal}(\mathbf{L}/\mathbf{F})$ in $W(\mathrm{E}_{6}).$ We remind the reader that $\Gamma$ lies in the centralizer of $\mathrm{Aut}(S),$ since any $\phi\in\mathrm{Aut}(S)$ is defined over $\mathbf{F}.$
Lemma 5.9.
Let $S$ be a smooth cubic surface over a field $\mathbf{F}$ such that all $27$ exceptional lines are defined over $\mathbf{F}.$ Let $G$ be a subgroup of order $5$ in $\mathrm{Aut}(S).$ Then there is a quintuple of skew lines in quadratic position such that they form an orbit under the action of $G.$
Proof.
By Lemma 4.3 there is only one conjugacy class of order $5$ in $W(\mathrm{E}_{6}).$ Let us consider the element $g\in W(\mathrm{E}_{6})$ of order $5$ such that it acts on the indices of exceptional lines as the permutation $(23456).$ In particular, it permutes the lines
$L_{12},\;L_{13},\;L_{14},\;L_{15},\;L_{16}.$
By Lemma 5.8 these skew lines are in quadratic position. As up to conjugation the group $G$ is generated by $g,$ we get the desired.
Corollary 5.10.
Let $S$ be a smooth cubic surface over a finite field $\mathbf{F}$ such that all $27$ exceptional lines are defined over $\mathbf{F}.$ Let $G$ be a subgroup of order $5$ in $\mathrm{Aut}(S).$ Then there is a $\mathbb{Z}/5\mathbb{Z}$-equivariant blowdown $\pi:S\to\mathbb{P}^{1}\times\mathbb{P}^{1}.$
Proof.
By Lemma 5.9 there are $5$ skew lines in quadratic position which form a $G$-orbit. Therefore, blowing them down gives us a $\mathbb{Z}/5\mathbb{Z}$-equivariant map
$\pi:S\to X,$
where $X_{\overline{\mathbf{F}}}\simeq\mathbb{P}^{1}\times\mathbb{P}^{1}.$ Since all $27$ exceptional lines are defined on $S,$ we have the equality $\mathrm{rk}\mathrm{Pic}(S)=7,$ and, therefore,
$\mathrm{rk}\mathrm{Pic}(X)=\mathrm{rk}\mathrm{Pic}(S)-5=2.$
By Lemma 3.2 we get that $X\simeq\mathbb{P}^{1}\times\mathbb{P}^{1}.$
Lemma 5.11.
Let $S$ be a smooth cubic surface over $\mathbf{F}$ such that the image $\Gamma$ of the absolute Galois group $\mathrm{Gal}(\mathbf{F}^{sep}/\mathbf{F})$ in $W(\mathrm{E}_{6})$ is isomorphic to $\mathbb{Z}/2\mathbb{Z}.$ Let $G$ be a subgroup of order $5$ in $\mathrm{Aut}(S).$ Then there is a quintuple of skew lines in quadratic position which are defined over $\mathbf{F}$ such that they form an orbit under the action of $G.$
Proof.
By Corollary 4.6 the image of the Galois group is generated by an element lying in the conjugacy class $A_{1}.$ So up to conjugation this element is equal to $\sigma$ defined in Remark 5.5. The generator $\gamma$ of $\Gamma$ and the generator $g$ of $G$ are commute. As by Lemma 4.3 there is a unique element of order $10$ up to conjugation in $W(\mathrm{E}_{6})$, we get that the element $\gamma g$ is unique up to conjugation in $W(\mathrm{E}_{6}).$ Therefore, without loss of generality we can assume that $r=\gamma g=\sigma\cdot(23456).$ Thus, $\gamma=r^{5}$ and $g=r^{2}.$
By the definition of $\sigma$ all the lines
$L_{12},\;L_{13},\;L_{14},\;L_{15},\;L_{16}$
are defined over $\mathbf{F}.$ By Lemma 5.8 they are in general position. As up to conjugation the group $G$ is generated by $g,$ we get the desired.
Corollary 5.12.
Let $S$ be a smooth cubic surface over a finite field $\mathbf{F}$ such that the image $\Gamma$ of the absolute Galois group $\mathrm{Gal}(\overline{\mathbf{F}}/\mathbf{F})$ in $W(\mathrm{E}_{6})$ is isomorphic to $\mathbb{Z}/2\mathbb{Z}.$ Let $G$ be a subgroup of order $5$ in $\mathrm{Aut}(S).$ Then there is a $\mathbb{Z}/5\mathbb{Z}$-equivariant blowdown
$\pi:S\to Q,$
where $Q\simeq\mathrm{R}_{\mathbf{L}/\mathbf{F}}(\mathbb{P}^{1})$ such that $\mathbf{F}\subset\mathbf{L}$ is a separable quadratic extension of $\mathbf{F}.$
Proof.
By Lemma 5.11 there are $5$ skew lines in quadratic position such that they form a $G$-orbit (in particular, all of them are defined over $\mathbf{F}$). Therefore, blowing them down gives us a $\mathbb{Z}/5\mathbb{Z}$-equivariant map
$\pi:S\to X,$
where $X_{\overline{\mathbf{F}}}\simeq\mathbb{P}^{1}\times\mathbb{P}^{1}.$ As the group $\Gamma$ is not trivial, we obtain $\mathrm{rk}\mathrm{Pic}(S)\leqslant 6,$ and, therefore,
$\mathrm{rk}\mathrm{Pic}(X)=\mathrm{rk}\mathrm{Pic}(S)-5=1.$
By Lemma 3.2 we get that $X\simeq Q,$ where $Q\simeq\mathrm{R}_{\mathbf{L}/\mathbf{F}}(\mathbb{P}^{1})$ such that $\mathbf{F}\subset\mathbf{L}$ is a separable quadratic extension of $\mathbf{F}.$
Now we give a simple lemma about the triviality of the image of the absolute Galois group in $W(\mathrm{E}_{6})$ for smooth cubic surface with large automorphism group.
Lemma 5.13.
Let $S$ be a smooth cubic surface over a field $\mathbf{F}$ such that its automorphism group is isomorphic either to $\mathrm{PSU}_{4}(\mathbb{F}_{2}),$ or to $\left(\mathbb{Z}/2\mathbb{Z}\right)^{4}\rtimes\mathfrak{A}_{5}.$ Then the image $\Gamma$ of the absolute Galois group $\mathrm{Gal}(\mathbf{F}^{sep}/\mathbf{F})$ in $W(\mathrm{E}_{6})$ is trivial.
Proof.
This directly follows from Lemma 4.9 since $\Gamma$ commutes with the automorphism group of $S.$
Finally, we prove two lemmas about the image of the absolute Galois group in the Weyl group $W(\mathrm{E}_{6})$ over $\mathbb{F}_{2^{2k+1}}.$
Lemma 5.14.
Let $S$ be a smooth cubic surface over $\mathbb{F}_{2^{2k+1}}$ such that $\mathrm{Aut}(S)$ contains a subgroup of order $5$ Then the image of the absolute Galois group in $W(\mathrm{E}_{6})$ is non-trivial.
Proof.
Assume the contrary. Then all exceptional lines on $S$ are defined over $\mathbb{F}_{2^{2k+1}}.$ By Lemma 5.9 there is $\mathbb{Z}/5\mathbb{Z}$-equivariant blowdown to $\mathbb{P}^{1}\times\mathbb{P}^{1}.$ However, by Lemma 2.3 the order of the group $\mathrm{Aut}(\mathbb{P}^{1}\times\mathbb{P}^{1})$ is not divisible by $5$
Lemma 5.15.
Let $S$ be a smooth cubic surface over $\mathbb{F}_{2^{2k+1}}$ such that $\mathrm{Aut}(S)\simeq\mathfrak{S}_{6}.$ Then the image $\Gamma$ of the absolute Galois group $\mathrm{Gal}\left(\overline{\mathbb{F}}_{2}/\mathbb{F}_{2^{2k+1}}\right)$ in $W(\mathrm{E}_{6})$ is isomorphic to $\mathbb{Z}/2\mathbb{Z}$ and generated by an element lying in the conjugacy class $A_{1}.$
Proof.
Let us consider an element $g\in\mathfrak{S}_{6}$ of order $5$ in $W(\mathrm{E}_{6}).$ By Lemma 4.4 its centralizer in $W(\mathrm{E}_{6})$ is isomorphic to $\mathbb{Z}/10\mathbb{Z}.$ The element $g$ does not lie in the centralizer of $\mathfrak{S}_{6}\subset W(\mathrm{E}_{6}),$ because it does not lie in the centre of $\mathfrak{S}_{6}.$ Therefore, the centralizer of $\mathfrak{S}_{6}$ in $W(\mathrm{E}_{6})$ lies in $\mathbb{Z}/2\mathbb{Z}.$ This means that $\Gamma\subseteq\mathbb{Z}/2\mathbb{Z}.$ By Lemma 5.14 the group $\Gamma$ is non-trivial. Therefore, the group $\Gamma$ is isomorphic to $\mathbb{Z}/2\mathbb{Z}.$
By Corollary 4.6 the group $\Gamma$ is generated by an element lying in the conjugacy class $A_{1}.$
6. Proof of the main results
In this section we study automorphism groups of smooth cubic surfaces over all finite fields of characteristic $2$ and prove Theorems 1.4 and 1.5. First of all, we prove the following lemmas about the automorphism group of maximal order of a smooth cubic surface over a finite field of characteristic $2$
Lemma 6.1.
[…]
Proof.
By Lemma 5.2 the automorphism group of a smooth cubic surface lies in $\mathrm{PSU}_{4}(\mathbb{F}_{2}).$
Let us consider smooth cubic surfaces over $\mathbb{F}_{4^{k}}.$ By Theorem 1.2 the automorphism group of the Fermat cubic surface over $\mathbb{F}_{4}$ is isomorphic to $\mathrm{PSU}_{4}(\mathbb{F}_{2}).$ This means that the automorphism group of maximal order of a smooth cubic surface over $\mathbb{F}_{4^{k}}$ is isomorphic to $\mathrm{PSU}_{4}(\mathbb{F}_{2}).$
Lemma 6.2.
Let $S$ be a smooth cubic surface over $\mathbb{F}_{2^{2k+1}}.$ Then $|\mathrm{Aut}(S)|\leqslant 720$ and in the case of equality one has $\mathrm{Aut}(S)\simeq\mathfrak{S}_{6}.$
Proof.
Let $S$ be a smooth cubic surface over $\mathbb{F}_{2^{2k+1}}$ such that its automorphism group is of maximal order among smooth cubic surfaces over $\mathbb{F}_{2^{2k+1}}.$ Then by Theorem 1.3 we have $|\mathrm{Aut}(S)|\geqslant 720.$ By Lemmas 4.7 and 5.2 we get $3$ candidates for being the largest automorphism group of a smooth cubic surface $S$:
$\mathfrak{S}_{6},\;\mathrm{PSU}_{4}(\mathbb{F}_{2})\;\text{and}\;\left(\mathbb{Z}/2\mathbb{Z}\right)^{4}\rtimes\mathfrak{A}_{5}.$
The groups $\mathrm{PSU}_{4}(\mathbb{F}_{2})$ and $\left(\mathbb{Z}/2\mathbb{Z}\right)^{4}\rtimes\mathfrak{A}_{5}$ are not automorphism groups of a smooth cubic surface over $\mathbb{F}_{2^{2k+1}},$ because otherwise by Lemma 5.13 the image of the absolute Galois group in $W(\mathrm{E}_{6})$ is trivial. However, by Lemma 5.14 it is impossible. This means that the largest automorphism group is isomorphic to $\mathfrak{S}_{6}.$
Now we are ready to prove our main results. Let us start with Theorem 1.5.
Proof of Theorem 1.5.
Let us prove assertion (i). Since $\Gamma$ is trivial, all lines on $S_{\overline{\mathbf{F}}}$ are defined over $\mathbf{F}.$ By Lemma 5.9 there are $5$ skew lines in quadratic position such that they form an orbit under the action of $\mathbb{Z}/5\mathbb{Z}\subset\mathrm{Aut}(S).$ By Corollary 5.10 we can blow them down $\mathbb{Z}/5\mathbb{Z}$-equivariantly and get $\mathbb{P}^{1}\times\mathbb{P}^{1}.$ The blowdown gives $5$ points in general position on $\mathbb{P}^{1}\times\mathbb{P}^{1}$ by Remark 5.7. By Propositions 3.5 and 3.6 we get that the set of these $5$ points is unique up to conjugation. Therefore, a smooth cubic surface such that its automorphism group contains $\mathbb{Z}/5\mathbb{Z}$ and $\Gamma$ is trivial is unique up to isomorphism. By Theorem 1.2 we get that the automorphism group of the Fermat cubic surface contains $\mathbb{Z}/5\mathbb{Z}.$ Thus, $S$ is isomorphic to the Fermat cubic surface.
Now let us prove assertion (ii). By Lemma 5.11 there are $5$ skew lines in quadratic position such that they form an orbit under the action of $\mathbb{Z}/5\mathbb{Z}\subset\mathrm{Aut}(S).$ By Corollary 5.12 we can blow them down $\mathbb{Z}/5\mathbb{Z}$-equivariantly and get the Weil restriction of scalars $Q\simeq\mathrm{R}_{\mathbf{L}/\mathbf{F}}(\mathbb{P}^{1}),$ where $\mathbf{F}\subset\mathbf{L}$ is a quadratic extension of $\mathbf{F}.$ The blowdown gives us $5$ points in general position on $Q$ by Remark 5.7. By Proposition 3.7 we get that the set of these $5$ points is unique up to automorphism. Therefore, a smooth cubic surface such that its automorphism group contains $\mathbb{Z}/5\mathbb{Z}$ and $\Gamma\simeq\mathbb{Z}/2\mathbb{Z}$ is unique up to isomorphism. By Theorem 1.3 we get that the automorphism group of the cubic surface (1.1) contains $\mathbb{Z}/5\mathbb{Z}.$ Thus, $S$ is isomorphic to the cubic surface (1.1).
Proof of Theorem 1.4.
First of all, let us consider smooth cubic surfaces over $\mathbb{F}_{4^{k}}.$ By Lemma 6.1 the maximal order automorphism group of such a surface is isomorphic to $\mathrm{PSU}_{4}(\mathbb{F}_{2}).$ It remains to show that a smooth cubic surface with such automorphism group is unique up to isomorphism. As $\mathrm{PSU}_{4}(\mathbb{F}_{2})$ contains $\mathbb{Z}/5\mathbb{Z}$ and by Lemma 5.13 the image of the absolute Galois group in $W(\mathrm{E}_{6})$ is trivial, we can apply Theorem 1.5(i) and get the desired.
Now let us consider smooth cubic surfaces over $\mathbb{F}_{2^{2k+1}}.$ By Lemma 6.2 the automorphism group of maximal order of such a surface is isomorphic to $\mathfrak{S}_{6}.$ It remains to show that a smooth cubic surface with such automorphism group is unique up to isomorphism. The group $\mathfrak{S}_{6}$ contains $\mathbb{Z}/5\mathbb{Z}.$ By Lemma 5.15 the image of the absolute Galois group in $W(\mathrm{E}_{6})$ is isomorphic to $\mathbb{Z}/2\mathbb{Z}.$ So we can apply Theorem 1.5(ii) and get the desired.
7. Fermat cubic surface over fields of characteristic $2$
In this section we study automorphism group of the Fermat cubic surface over fields of characteristic $2$ and prove Proposition 1.6 and Corollary 1.7.
Proof of Proposition 1.6.
First of all, note that we have
(7.1) $\mathbb{Z}/2\mathbb{Z}\times\mathfrak{S}_{4}\subseteq\mathrm{Aut}(\mathcal{F}),$
where $\mathfrak{S}_{4}$ permutes the coordinates and $\mathbb{Z}/2\mathbb{Z}$ is generated by the element
(7.2) $h=\begin{pmatrix}0&1&1&1\\ 1&0&1&1\\ 1&1&0&1\\ 1&1&1&0\end{pmatrix}.$
From the equations of lines on the Fermat cubic surface we get that the Galois group of the field of definition of all lines on $\mathcal{F}$ is isomorphic to $\mathbb{Z}/2\mathbb{Z}.$ The generator $g$ of this group maps $\omega$ to $\omega^{2},$ where $\omega\in\overline{\mathbf{F}}$ is a non-trivial cube root of unity. So again from the equations of lines on the Fermat cubic surface one can see that the element $g$ fixes $3$ lines, and interchanges the lines in the six pairs of intersecting lines and in the six pairs of non-intersecting lines. By this description of the element from Lemma 5.3 we see that the element $g$ in the Weyl group $W(\mathrm{E}_{6})$ lies in the conjugacy class $A_{1}^{3}.$ By Lemma 4.3 the order of its centralizer is $96$ By the same lemma the determinant of $g$ is equal to $-1.$ As both $g$ and $\mathrm{Aut}(\mathcal{F})$ lie in the centralizer of $g$ and the element $g$ does not lie in $\mathrm{Aut}(\mathcal{F}),$ we have the isomorphism in (7.1) by the comparison of the orders.
Proof of Corollary 1.7.
By Theorem 1.4(ii) the automorphism group of the cubic (1.1) is isomorphic to $\mathfrak{S}_{6},$ while by Proposition 1.6 the automorphism of the Fermat cubic surface is isomorphic to $\mathbb{Z}/2\mathbb{Z}\times\mathfrak{S}_{4}.$ This proves assertion (i).
The isomorphism between (1.1) and the Fermat cubic surface over $\mathbb{F}_{4}$ is given by the collineation
$[x:y:z:t]\mapsto[x+y:\omega x+\omega^{2}y:z+t:\omega z+\omega^{2}t],$
where $\omega\in\mathbb{F}_{4}$ is a non-trivial cube root of unity, which maps the cubic surface (1.1) to the Fermat cubic surface. This proves assertion (ii).